OpenStudy (anonymous):

*help requested* *fan and medal awarded* Simplify 6(3u - 2). 9u - 12 18u - 12 18u + 12 18u - 2 Evaluate: -4 - (-6) 2 -10 -2 24 If A = {negative even integers} and B = {negative odd integers}, what is A union B? {0} {negative integers} {integers} { }

OpenStudy (anonymous):

1) Distributive property. a(b + c) = ab + ac 6(3u - 2) = 18u - 12

OpenStudy (anonymous):

2) Two negatives equal a positive. So... -4 - (-6) -4 + 6 Can you solve that^

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

No. Try looking at a number line. 

OpenStudy (anonymous):

i got to get off the pc so can u jus explain the rest and ill fan and give umedal when im back ?

OpenStudy (anonymous):

*u medal

OpenStudy (anonymous):

Add 6